The width of Matt's new rectangular vegetable garden is 4 feet less than the length. If the perimeter of the garden is 56 feet, find the dimensions of the garden.
Answer by ewatrrr(24785)

 
Hi,

P = 2L + 2w = 56 ft 0r L + w = 28ft
Question States*** w =+L-4
L + w = 28ft
L + (L-4) = 28ft
2L = 32
L = 16ft
garden is 12ft by 16ft